Advertisements
Advertisements
प्रश्न
The odds against student X solving a business statistics problem are 8: 6 and odds in favour of student Y solving the same problem are 14: 16 What is the probability that neither solves the problem?
Advertisements
उत्तर
Let A be the event that X solves the problem B be the event that Y solves the problem.
Since the odds against student X solving the problem are 8: 6
∴ Probability of occurrence of event A is given by.
P(A) = `6/(8 + 6) = 6/14` and
P(A') = 1 – P(A) = `1 - 6/14 = 8/14`
Also, the odds in favour of student Y solving the problem are 14: 16
∴ Probability of occurrence of event B is given by
P(B) = `14/(14 + 16) = 14/30` and
P(B') = 1 – P(B) = `1 - 14/30 = 16/30`
Now A and B are independent events.
∴ A' and B' are independent events.
P (neither solves the problem) = P(A' ∩ B')
= P(A') . P(B')
= `8/14 xx 16/30`
= `32/105`
APPEARS IN
संबंधित प्रश्न
Probability of solving specific problem independently by A and B are `1/2` and `1/3` respectively. If both try to solve the problem independently, find the probability that
- the problem is solved
- exactly one of them solves the problem.
A problem in statistics is given to three students A, B, and C. Their chances of solving the problem are `1/3`, `1/4`, and `1/5` respectively. If all of them try independently, what is the probability that, exactly two students solve the problem?
The odds against a certain event are 5: 2 and odds in favour of another independent event are 6: 5. Find the chance that at least one of the events will happen.
The odds against student X solving a business statistics problem are 8: 6 and odds in favour of student Y solving the same problem are 14: 16 What is the chance that the problem will be solved, if they try independently?
A family has two children. Find the probability that both the children are girls, given that atleast one of them is a girl.
Solve the following:
Find the probability that a year selected will have 53 Wednesdays
If A and B are independent events such that 0 < P(A) < 1 and 0 < P(B) < 1, then which of the following is not correct?
For a loaded die, the probabilities of outcomes are given as under:
P(1) = P(2) = 0.2, P(3) = P(5) = P(6) = 0.1 and P(4) = 0.3. The die is thrown two times. Let A and B be the events, ‘same number each time’, and ‘a total score is 10 or more’, respectively. Determine whether or not A and B are independent.
Let E1 and E2 be two independent events such that P(E1) = P1 and P(E2) = P2. Describe in words of the events whose probabilities are: 1 – (1 – P1)(1 – P2)
Two dice are tossed. Find whether the following two events A and B are independent: A = {(x, y): x + y = 11} B = {(x, y): x ≠ 5} where (x, y) denotes a typical sample point.
Let A and B be two events such that P(A) = `3/8`, P(B) = `5/8` and P(A ∪ B) = `3/4`. Then P(A|B).P(A′|B) is equal to ______.
If A and B are independent, then P(exactly one of A, B occurs) = P(A)P(B') + P(B)P(A')
Let A and B be two events. If P(A | B) = P(A), then A is ______ of B.
Two events 'A' and 'B' are said to be independent if
Let EC denote the complement of an event E. Let E1, E2 and E3 be any pairwise independent events with P(E1) > 0 and P(E1 ∩ E2 ∩ E3) = 0. Then `"P"(("E"_2^"C" ∩ "E"_3^"C")/"E"_1)` is equal to ______.
The probability of the event A occurring is `1/3` and of the event B occurring is `1/2`. If A and B are independent events, then find the probability of neither A nor B occurring.
Which is the equivalent and most commonly used test for independent events E and F?
A die is thrown once. If F is the event that the number appearing is even, which set represents F?
For \[F=\{2,4,6\}\] when a die is thrown once, what is \[P(F)\]?
